OpenStudy (anonymous):

(a+5)/(a^2-9)-(a+2)/(a^2-6a+9)

OpenStudy (dumbcow):

\[\frac{a+5}{(a+3)(a-3)}-\frac{a+2}{(a-3)^{2}} = \frac{(a+5)(a-3) - (a+2)(a+3)}{(a+3)(a-3)^{2}}\]

OpenStudy (anonymous):

\[-(3 (a+7))/((a-3)^2 (a+3)) \]